# Quark-Gluon Plasma, Charmonia, Ultra-relativistic Heavy-Ion Collisions

### v1

@@ 2.3 Charmonium equilibrium limits
2.4 Initial Conditions and Medium Evolution
Refer to captionRefer to caption
Figure 4: Our parameterizations of the suppression of the $c\bar{c}$ production cross section due to nuclear shadowing of the parton distribution functions, as a function of $p_{\rm T}$ (left panel) in p-Pb collisions in terms of the product at forward and backward rapidities and compared to ALICE data Adam et al. (2015), and versus $N_{\rm part}$ in Pb-Pb collisions (right panel).
The cross section for $c\bar{c}$ pair production in pp collisions is a key input for the equilibrium limits and thus controls the amount of regeneration. It is usually applied for a specific rapidity interval as $N_{c\bar{c}}=\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}yN_{\rm coll}$, where $N_{\rm coll}$ denotes the number of binary nucleon-nucleon collisions at a given collision energy and centrality (estimated from the optical Glauber model). Similarly, we obtain the initial number of charmonium states, which are required to determine the initial condition for the rate equation. In Table 1 we summarize the charm/onium cross sections for pp collisions at 5.02 TeV and the ratio of $\psi(2S)$ to $J/\psi$. The open-charm cross section at forward rapidity has been deduced from its counterpart at midrapidity through the rapidity dependence outlined in Ref. Bierlich et al. (2023). This extrapolation yields a value $0.72\pm 0.07$, where we incorporate an uncertainty of ±$\pm 10\%$.
Table 1: Charm/onium cross sections for pp collisions at 5.02 TeV and the ratio for $\psi(2S)$ to $J/\psi$.
@@ 2.4 Initial Conditions and Medium Evolution
$\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}y$ (mb) $1.165\pm 0.133$ Acharya et al. (2022) $0.72\pm 0.07$
$\mathrm{d}\sigma_{J/\psi}/\mathrm{d}y$ ($\mu$b) $5.64$ Acharya et al. (2019) $3.93$ Acharya et al. (2023)
Direct-production cross section ratio in pp collisions
$N_{\psi(2S)}^{pp}/N_{J/\psi}^{pp}$ $0.147$ Acharya et al. (2023)
A modification of the charm(onium) cross section due to cold-nuclear-matter effects, commonly referred to as nuclear shadowing, is estimated using ALICE data Adam et al. (2015) on $J/\psi$ production in p-Pb collisions at forward and backward rapidities, $42.5<\left|y\right|<$ 4. We fit the product of the measured forward and backward nuclear modification factors, $R_{\rm pPb}$, which can be interpreted as the net effect of shadowing in a Pb-Pb collision, as shown in the left panel of Figure 4. In earlier applications of our transport approach to p-Pb collisions at the LHC Du and Rapp (2019), where a short-lived QGP is predicted to be formed, it was found that about 10-20% of the inclusive $J/\psi$ suppression is from hot-matter effects (primarily from feeddown of suppressed excited states), which must be “corrected" for when assessing the shadowing effect. Therefore, we adopt a baseline of 10-30% suppression of the integrated yield stemming from shadowing, with a $p_{\rm T}$ dependence that reproduces the forward-backward $R_{\rm pPb}$ product. The $N_{\rm part}$-dependence of the shadowing for the $c\bar{c}$ cross section is displayed in the right panel of Figure 4.
In our comparisons to experimental results for Pb-Pb collisions discussed below, we incorporate the experimental uncertainties of the charm-anticharm cross section $\sigma_{c\bar{c}}$, denoted by $\Delta\sigma_{c\bar{c}}$, along with the shadowing factor $S$ and its uncertainty, $\Delta S$. The uncertainty of the pp cross section and its shadowing is combined into an effective cross section, $\tilde{\sigma}_{c\bar{c}}$, as follows:
@@
Figure 5: Fraction of bottom-hadron decay feeddown to inclusive $J/\psi$ production as measured by ALICE, ATLAS and CMS in pp collisions at the LHC Andronic et al. (2016); our functional fit is shown by the solid line.
Data for inclusive $J/\psi$ production also include feeddown contributions from final-state decays. For the “prompt" feeddown from electromagnetic and strong decays of excited states, we account for 8% and 25% from $\psi(2S)$ and $\chi_{c}$ mesons, respectively Andronic et al. (2016). Additionally, there is also a “non-prompt" fraction from bottom decay feeddown (unless explicitly removed experimentally). We estimate this fraction from available data Andronic et al. (2016) in pp collisions at the LHC, cf. Figure 5. It amounts to around 5% at $p_{\rm T}$=0 and increases to approximately 50% at $p_{\rm T}$=20 GeV. We fit the data using the empirical parameterization
@@
The implementation of the rate equation into URHICs requires the space-time evolution of the volume and temperature of the expanding medium. Toward this end, we employ a rather simple fireball model for a cylindrical, isotropic, and isentropic evolution, as elaborated in previous studies Grandchamp and Rapp (2001); Zhao and Rapp (2010, 2011) for SPS, RHIC and LHC energies. The volume expansion, reminiscent of hydrodynamic models, essentially corresponds to a time-dependent blastwave model with a collective flow at thermal freezeout ($T_{\rm fo}\simeq$ 100 MeV in central Pb-Pb collisions) that reproduces measured $p_{T}$ spectra of light hadrons. The total entropy in the fireball, calculated from the observed multiplicity of charged particles using a hadron resonance gas model Rapp (2002), is assumed to remain constant throughout the adiabatic expansion. By monitoring the entropy density, $s(\tau)=S_{\rm tot}/V_{\rm FB}$, at each time $\tau$ we can infer the fireball temperature once we specify the equation of state (EoS) for the medium, i.e., $s(T)$. For the QGP phase, we adopt an ideal gas of massive quarks and gluons, while the hadronic phase is represented by a non-interacting gas of resonant states, including mesons and anti-/baryons with masses of up to 2 GeV. The critical temperature is set to $T_{c}$=180 MeV (we have checked that employing a more realistic EoS based on lattice-QCD data, with a continuous transition into a hadron resonance gas, has negligible impact on quarkonium kinetics Du et al. (2017)). Figure 6 illustrates the resultant temperature evolution over time across various centralities at LHC energies.
@@
Right after the near-instantaneous formation of a $c\bar{c}$ pair which would develop into a $\psi$ in a pp collision, we incorporate initial formation time effects to approximate the (quantum) expansion dynamics of evolving into a fully formed bound state. Contrary to expectations derived from a classical cross section perspective where the transverse size (i.e., cross sectional area) would grow quadratically with time, we utilize a scaling that is linear with time Farrar et al. (1988) and scale down the reaction rates by a factor of $\tau/\tau_{\rm form}$ for $\tau\leq\tau_{\rm form}$. The quantum formation times are estimated based on energy uncertainties associated with the splitting in binding energies, i.e., $\tau_{\mathrm{form}}(J/\psi,\psi(2S),\chi_{c})=1,2,2\,\text{fm}/c$, respectively.
@@ 3 Time and Centrality Dependence of Charmonium Yields at the LHC
where $N_{\psi}^{pp}$ denotes the inclusive $\psi$ yield in pp collisions at the same collision energy. The number of nucleon participants, $N_{\rm part}$, is estimated from a Glauber model for a given impact parameter, $b$, and serves as a measure of the centrality of the nuclear collision. Unless otherwise stated, the denominator of the $R_{\rm AA}$ will include both prompt and non-prompt feeddown contributions, and the reaction rates are for quasifree dissociation.
@@ 3.1 Time evolution of charmonium yields
The time evolution of the nuclear-modification factors for $J/\psi$ and $\psi(2S)$ in 0-20% and 60%-90% central Pb-Pb (5.02 TeV) collisions, including contributions from both primordial and regenerated yields, alongside their equilibrium limits, is shown in Figure 7, where the numerator only includes direct production (i.e., excluding feeddown). In the initial phases of the QGP evolution in central collisions, both primordial charmonium states undergo strong suppression, with the $\psi(2S)$ yield essentially being wiped out. In contrast, peripheral collisions show significantly less suppression, especially for the $J/\psi$, due to a lower fireball temperature and shorter lifetime. Nevertheless, a marked suppression of the $\psi(2S)$ is still operative as its reaction rates are still appreciable at the critical temperature as well as in the hadronic phase.
Figure 7: Time evolution of charmonium kinetics in central (left panel) and peripheral (right panel) Pb-Pb collisions at the LHC at forward rapidity. Blue and orange lines represent the direct $J/\psi$ and $\psi(2S)$ yields, respectively, where the solid and dashed lines correspond to the suppressed primordial and regenerated contributions, respectively, while the dashed-dotted lines are the pertinent equilibrium limits (including the thermal relaxation time correction). The calculations are carried out with the charm cross section $\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}y=0.72\,$mb including an up to 20% shadowing in central collisions (which is down to about 4% at 60-90% centrality).
@@
Figure 8: Centrality dependence of charmonium $R_{\rm AA}$’s in Pb-Pb(5.02 TeV) collisions at the LHC. The upper (lower) panels are for mid- (forward) rapidities and the left (right) panels are for $J/\psi$ ($\psi(2S)$). The bands for the primordial (orange), regenerated (red), and total (blue) components include uncertainties from the initial charm cross section and the shadowing effect (added in quadrature). The experimental cuts in pair $p_{\rm T}$ serve to suppress the contribution from coherent photoproduction and are also applied to our calculations based on the $p_{\rm T}$ spectra computed in Sec. 4. The calculations are compared to ALICE $J/\psi$ data from 2018 (brown) and 2023 (black) Adam et al. (2017); Bai (2021); Acharya et al. (2024), and $\psi(2S)$ data Acharya et al. (2024) data.
@@ 3.2 Centrality dependence
The centrality dependence of $J/\psi$ and $\psi(2S)$ yields is obtained by evaluating Eq. 1 with initial conditions determined for a given (average) participant number, complemented with prompt and non-prompt feeddown contributions. The results for inclusive $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, as a function of $N_{\rm part}$ at mid- and forward rapidity, are summarized in Figure 8. The $J/\psi$ yield exhibits the well-established behavior at LHC energies: a strong initial suppression that sets in rather gradually with centrality and is taken over by regeneration contributions at participant numbers of around 100-150. At mid-rapidity, both suppression and regeneration are slightly stronger than at forward rapidity, due to a hotter medium and a larger charm cross section, respectively. The interplay of these mechanisms produces a fairly flat centrality dependence for the total $R_{\rm AA}$, with a mild rise at mid-rapidity, again due to the larger charm production. Overall, the ALICE $J/\psi$ data are reasonably well described Adam et al. (2017); Acharya et al. (2024, 2024), with a preference for shadowing on the weaker side of our central values. For the $\psi(2S)$, the right panels in Figure 8 show our predictions based on Refs. Zhao and Rapp (2011); Du and Rapp (2015) with the updated inputs as discussed in the previous section. Compared to the $J/\psi$, a much steeper suppression of the initial production is found due to the larger rates and smaller dissociation temperatures, leading to a near-complete suppression for $N_{\rm part}\gtrsim$ 100. Detailed balance causes a regeneration contribution that takes over from the primordial yield at $N_{\rm part}$ as low as $\sim$40. The resulting inclusive $R_{\rm AA}$ is also quite flat, but leveling off at a substantially smaller value than for the $J/\psi$, chiefly due to the smaller equilibrium limit caused by its larger mass.
@@
Figure 9: The ratio of $\psi(2S)$ over $J/\psi$ as a function of $N_{\rm part}$ in Pb-Pb(5.02 TeV) collisions at forward rapidity, compared to ALICE data Acharya et al. (2024). The bands indicate the uncertainty of the $\psi(2S)$ dissociation temperature around the mixed phase, $T_{\rm diss}=179-180$ MeV.
@@
We also evaluate the ratio of $\psi(2S)$ to $J/\psi$, which mitigates some of the uncertainties in the individual yields, such as the input charm cross section. To illustrate the uncertainty caused by the assumption of the $\psi(2S)$ dissociation temperature (below which regeneration sets in), we vary it in the range of 179-180 MeV (at the beginning of the mixed phase, which is our default) as a lower and upper limit, respectively. Using the pertinent branching fractions for dilepton decays of $\text{BR}(\psi(2S)\rightarrow\mu^{+}\mu^{-})=8\times 10^{-3}$ and $\text{BR}(J/\psi\rightarrow\mu^{+}\mu^{-})=5.961\times 10^{-2}$ Zyla et al. (2020), the predicted $N_{\rm part}$ dependence is shown in Figure 9. After an initial rather sharp drop in peripheral collisions (driven by the strong suppression of primordial $\psi(2S)$’s) the ratio essentially saturates and turns out to be consistent with the experimental findings reported by the ALICE collaboration Acharya et al. (2024). Our predictions are significantly larger than the results from the statistical hadronization model Andronic et al. (2019, 2018), which level off at approximately $0.05$ in central collisions. As discussed above, the reason for this is that in our transport approach the regeneration of $J/\psi$’s does not reach its equilibrium limit, while the $\psi(2S)$ does, albeit at lower temperatures, recall Figure 7.
@@
Next, we return to scrutinizing the impact of the gluo-dissociation processes on our results by incorporating the pertinent rates into the rate equation [^1] Strictly speaking, this is not following the philosophy of our approach where we adjust the main parameters, i.e., the effective coupling constant, $\alpha_{s}$, in the quasifree rates as well as the thermal relaxation rate of charm quarks, to match SPS and RHIC data Grandchamp and Rapp (2001); Zhao and Rapp (2010). However, the small impact of gluo-dissociation on our results renders this exercise rather obsolete.. In Figure 10, we present results for the time evolution in central collisions (left panel; for direct production) and the centrality dependence (right panel; for inclusive production) of the regenerated, primordial and total $J/\psi$ $R_{\rm AA}$ with and without gluo-dissociation at forward rapidity. Note that a higher rate implies both stronger suppression and increased regeneration. For peripheral and semi-central collisions these two effects essentially compensate each other, while for central collisions, the total $J/\psi$ yield increases by $\sim$6due to the enhanced regeneration.
Refer to captionRefer to caption
@@
Figure 10: The time evolution (left panel) and centrality dependence (right panel, additionally including $B$ feeddown) of the regenerated (red), primordial (orange), and total (blue) $J/\psi$ production, with the bands illustrating the uncertainty as to whether gluo-dissociation rates are accounted for or not. The same values for shadowing and the $c\bar{c}$ cross section as in Figure 7 are used, and the ALICE data are from Ref. Acharya et al. (2024).
@@
Let us also briefly come back to the effect of the updated binding energies as compared to our previous calculations, as mentioned in the first paragraph of Sec. 2.2. The impact of the somewhat larger binding energies (most notably around $T$$\simeq$220 MeV) on regeneration is negligible, as one finds an almost complete compensation between the reduced rates and the larger equilibrium limits. The smaller rates do, however, affect the primordial suppression. The maximal effect for central Pb-Pb collisions amounts to an increase of the primordial contribution by about 0.03 units in terms of the $R_{\rm AA}$.

### v2

@@ 2.4 Initial Conditions and Medium Evolution
$\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}y$ (mb) $1.165\pm 0.133$ ALICE:2021dhb; ALICE:2023sgl $0.72\pm 0.07$
[…]
Direct-production cross section ratio in $p$ collisions
$N_{\psi(2S)}^{pp}/N_{J/\psi}^{pp}$ $0.147$ ALICE:2021qlw
Figure 4: Our parameterizations of the suppression of the $c\bar{c}$ production cross section due to nuclear shadowing of the parton distribution functions, as a function of $p_{\rm T}$ (left panel) in p-Pb collisions in terms of the product at forward and backward rapidities and compared to ALICE data ALICE:2015sru, and versus $N_{\rm part}$ in Pb-Pb collisions (right panel).
The cross section for $c\bar{c}$ pair production in $p$ collisions is a key input for the equilibrium limits and thus controls the amount of regeneration. It is usually applied for a specific rapidity interval as $N_{c\bar{c}}=\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}yN_{\rm coll}$, where $N_{\rm coll}$ denotes the number of binary nucleon-nucleon collisions at a given collision energy and centrality (estimated from the optical Glauber model Miller:2007ri; for the partonic production processes under consideration here, which are mostly driven by gluon fusion, no distinction is made for the proton-neutron and neutron-neutron collisions). Similarly, we obtain the initial number of charmonium states, which are required to determine the initial condition for the rate equation. In Table 1 we summarize the charm/onium cross sections for $p$ collisions at 5.02 TeV and the ratio of $\psi(2S)$ to $J/\psi$ (we note that our values for the charm cross section, taken from Ref. ALICE:2021dhb, are well in line with the most recent assessment in Ref. ALICE:2023sgl). The open-charm cross section at forward rapidity has been deduced from its counterpart at midrapidity through the rapidity dependence outlined in Ref. Bierlich:2023ewv. This extrapolation yields a value $0.72\pm 0.07$, where we incorporate an uncertainty of $\pm 10\%$.
A modification of the charm(onium) cross section due to cold-nuclear-matter effects, commonly referred to as nuclear shadowing, is estimated using ALICE data ALICE:2015sru on $J/\psi$ production in p-Pb collisions at forward and backward rapidities, $2.5<\left|y\right|<4$. We fit the product of the measured forward and backward nuclear modification factors, $R_{\rm pPb}$, which can be interpreted as the net effect of shadowing in a Pb-Pb collision, as shown in the left panel of Figure 4. Our fit is not inconsistent with recent nuclear parton distribution functions, see, e.g., Ref. Kusina:2017gkz. In addition, we use the same parameterization at mid-rapidity, which is compatible with recent ALICE data as well ALICE:2021lmn. In earlier applications of our transport approach to p-Pb collisions at the LHC Du:2018wsj, where a short-lived QGP is predicted to be formed, it was found that about 10-20% of the inclusive $J/\psi$ suppression is from hot-matter effects (primarily from feeddown of suppressed excited states), which must be “corrected" for when assessing the shadowing effect. Therefore, we adopt a baseline of 10-30% suppression of the integrated yield stemming from shadowing, with a $p_{\rm T}$ dependence that reproduces the forward-backward $R_{\rm pPb}$ product. The $N_{\rm part}$-dependence of the shadowing for the $c\bar{c}$ cross section is displayed in the right panel of Figure 4.
In our comparisons to experimental results for Pb-Pb collisions discussed below, we incorporate the experimental uncertainties of the charm-anticharm cross section $\sigma_{c\bar{c}}$, denoted by $\Delta\sigma_{c\bar{c}}$, along with the shadowing factor $S$ and its uncertainty, $\Delta S$. The uncertainty of the $p$ cross section and its shadowing is combined into an effective cross section, $\tilde{\sigma}_{c\bar{c}}$, as follows:
@@
Figure 5: Fraction of bottom-hadron decay feeddown to inclusive $J/\psi$ production as measured by ALICE, ATLAS and CMS in $p$ collisions at the LHC Andronic:2015wma; our functional fit is shown by the solid line.
Data for inclusive $J/\psi$ production also include feeddown contributions from final-state decays. For the “prompt" feeddown from electromagnetic and strong decays of excited states, we account for 8% and 25% from $\psi(2S)$ and $\chi_{c}$ mesons, respectively Andronic:2015wma. Additionally, there is also a “non-prompt" fraction from bottom decay feeddown (unless explicitly removed experimentally). We estimate this fraction from available data Andronic:2015wma in $p$ collisions at the LHC, cf. Figure 5. It amounts to around 5% at $p_{\rm T}$=0 and increases to approximately 50% at $p_{\rm T}=20\,{\rm GeV}/c$. We fit the data using the empirical parameterization
@@
The implementation of the rate equation into URHICs requires the space-time evolution of the volume and temperature of the expanding medium. Toward this end, we employ a rather simple fireball model for a cylindrical, isotropic, and isentropic evolution, as elaborated in previous studies Grandchamp:2002wp; Zhao:2010nk; Zhao:2011cv for SPS, RHIC and LHC energies. The volume expansion, reminiscent of hydrodynamic models, essentially corresponds to a time-dependent blastwave model with a collective flow at thermal freezeout ($T_{\rm fo}\simeq$ 100 MeV in central Pb-Pb collisions) that reproduces measured $p_{T}$ spectra of light hadrons. The total entropy in the fireball, calculated from the observed multiplicity of charged particles using a hadron resonance gas model, is assumed to remain constant throughout the adiabatic expansion. By monitoring the entropy density, $s(\tau)=S_{\rm tot}/V_{\rm FB}$, at each time $\tau$ we can infer the fireball temperature once we specify the equation of state (EoS) for the medium, i.e., $s(T)$. For the QGP phase, we adopt an ideal gas of massive quarks and gluons, while the hadronic phase is represented by a non-interacting gas of resonant states, including mesons and anti-/baryons with masses of up to 2 GeV. The critical temperature is set to $T_{c}$=180 MeV (we have checked that employing a more realistic EoS based on lattice-QCD data, with a continuous transition into a hadron resonance gas, has negligible impact on quarkonium kinetics Du:2017qkv). While the assumption of an isotropic volume, $V_{\rm FB}$, without spatial temperature gradients is rather schematic, it enables a straightforward calculation of the charm-quark fugacity factor, $\gamma_{c}$ in Eq. (10). The latter governs the equilibrium limit, Eq. (9), of the different charmonium states and is therefore key in obtaining a reliable gain term for the regeneration contribution. The extension to individual cells with different temperatures, as figuring, e.g., in a hydrodynamic evolution, renders this more challenging, but will be addressed in future work. Figure 6 illustrates the resultant temperature evolution as a function of proper time across various centralities at LHC energies int he forward rapidity region. At central rapidities, where the observed charged-particle multiplicities are about 20% larger for the same centrality class, the initial temperatures increase by about 6%.
@@
Right after the near-instantaneous formation of a $c\bar{c}$ pair which would develop into a $\psi$ in a $p$ collision, we incorporate initial formation time effects to approximate the (quantum) expansion dynamics of evolving into a fully formed bound state. Contrary to expectations derived from a classical cross section perspective where the transverse size (i.e., cross sectional area) would grow quadratically with time, we utilize a scaling that is linear with time Farrar:1988me and scale down the reaction rates by a factor of $\tau/\tau_{\rm form}$ for $\tau\leq\tau_{\rm form}$. The quantum formation times are estimated based on energy uncertainties associated with the splitting in binding energies, i.e., $\tau_{\mathrm{form}}(J/\psi,\psi(2S),\chi_{c})=1,2,2\,\text{fm}/c$, respectively.
@@ 3 Time and Centrality Dependence of Charmonium Yields at the LHC
where $N_{\psi}^{pp}$ denotes the inclusive $\psi$ yield in $p$ collisions at the same collision energy. The number of nucleon participants, $N_{\rm part}$, is estimated from a Glauber model for a given impact parameter, $b$ Miller:2007ri, and serves as a measure of the centrality of the nuclear collision. Unless otherwise stated, the denominator of the $R_{\rm AA}$ will include both prompt and non-prompt feeddown contributions and utilize the central values for the input cross sections at given rapidity, and the reaction rates are for quasifree dissociation.
@@ 3.1 Time evolution of charmonium yields
Figure 7: Time evolution of charmonium kinetics in central (left panel) and peripheral (right panel) Pb-Pb collisions at the LHC at forward rapidity. Blue and orange lines represent the direct $J/\psi$ and $\psi(2S)$ yields, respectively, where the solid and dashed lines correspond to the suppressed primordial and regenerated contributions, respectively, while the dashed-dotted lines are the pertinent equilibrium limits (including the thermal relaxation time correction). The calculations are carried out a charm cross section of $\mathrm{d}\sigma_{c\bar{c}}/\mathrm{d}y=0.72\,$mb (the central value at forward rapidity, recall Table 1) including an up to 20% shadowing in central collisions (which is down to about 4% at 60-90% centrality).
To illustrate the time evolution of the nuclear-modification factors for $J/\psi$ and $\psi(2S)$, we focus on 0-20% and 60%-90% central Pb-Pb (5.02 TeV) collisions at forward rapidity, cf. Figure 7. We include contributions from both primordial and regenerated yields, alongside their equilibrium limits, where the numerator only includes direct production (i.e., excluding prompt of weak-decay feeddown). In the initial phases of the QGP evolution in central collisions, both primordial charmonium states undergo strong suppression, with the $\psi(2S)$ yield essentially being wiped out. In contrast, peripheral collisions show significantly less suppression, especially for the $J/\psi$, due to a lower fireball temperature and shorter lifetime. Nevertheless, a marked suppression of the $\psi(2S)$ is still operative as its reaction rates are still appreciable at the critical temperature as well as in the hadronic phase.
@@
Figure 8: Centrality dependence of charmonium $R_{\rm AA}$’s in Pb-Pb(5.02 TeV) collisions at the LHC. The upper (lower) panels are for mid- (forward) rapidities and the left (right) panels are for $J/\psi$ ($\psi(2S)$). The bands for the primordial (orange), regenerated (red), and total (blue) components include uncertainties from the initial charm cross section and the shadowing effect (added in quadrature). The experimental cuts in pair $p_{\rm T}$ serve to suppress the contribution from coherent photoproduction and are also applied to our calculations based on the $p_{\rm T}$ spectra computed in Sec. 4. The calculations are compared to ALICE $J/\psi$ data from 2018 (brown) and 2023 (black) ALICE:2016flj; Bai:2020svs; ALICE:2023gco, and $\psi(2S)$ data ALICE:2022jeh data.
@@ 3.2 Centrality dependence
The centrality dependence of $J/\psi$ and $\psi(2S)$ yields is obtained by evaluating Eq. 1 with initial conditions determined for a given (average) participant number, complemented with prompt and non-prompt feeddown contributions. The results for inclusive $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, as a function of $N_{\rm part}$ at mid- and forward rapidity, are summarized in Figure 8. The $J/\psi$ yield exhibits the well-established behavior at LHC energies: a strong initial suppression that sets in rather gradually with centrality and is taken over by regeneration contributions at participant numbers of around 100-150. At mid-rapidity, both suppression and regeneration are slightly stronger than at forward rapidity, due to a hotter medium and a larger charm cross section, respectively. The interplay of these mechanisms produces a fairly flat centrality dependence for the total $R_{\rm AA}$, with a mild rise at mid-rapidity, again due to the larger charm production. Overall, the ALICE $J/\psi$ data are reasonably well described ALICE:2016flj; ALICE:2022jeh; ALICE:2023gco, with a preference for shadowing on the weaker side of our central values. For the $\psi(2S)$, the right panels in Figure 8 show our predictions based on Refs. Zhao:2011cv; Du:2015wha with the updated inputs as discussed in the previous section. Compared to the $J/\psi$, a much steeper suppression of the initial production is found due to the larger rates and smaller dissociation temperatures, leading to a near-complete suppression for $N_{\rm part}\gtrsim$ 100. Detailed balance causes a regeneration contribution that takes over from the primordial yield at $N_{\rm part}$ as low as $\sim$40. The resulting inclusive $R_{\rm AA}$ is also quite flat, but leveling off at a substantially smaller value than for the $J/\psi$, chiefly due to the smaller equilibrium limit caused by its larger mass.
@@
Figure 9: The ratio of $\psi(2S)$ over $J/\psi$ as a function of $N_{\rm part}$ in Pb-Pb(5.02 TeV) collisions at forward rapidity, compared to ALICE data ALICE:2022jeh. The bands indicate the uncertainty of the $\psi(2S)$ dissociation temperature around the mixed phase, $T_{\rm diss}=179-180$ MeV.
@@
We also evaluate the ratio of $\psi(2S)$ to $J/\psi$, which mitigates some of the uncertainties in the individual yields, such as the input charm cross section. To illustrate the uncertainty caused by the assumption of the $\psi(2S)$ dissociation temperature (below which regeneration sets in), we vary it in the range of 179-180 MeV (at the beginning of the mixed phase, which is our default) as a lower and upper limit, respectively. Using the pertinent branching fractions for dilepton decays of $\text{BR}(\psi(2S)\rightarrow\mu^{+}\mu^{-})=8\times 10^{-3}$ and $\text{BR}(J/\psi\rightarrow\mu^{+}\mu^{-})=5.961\times 10^{-2}$ ParticleDataGroup:2020ssz, the predicted $N_{\rm part}$ dependence is shown in Figure 9. After an initial rather sharp drop in peripheral collisions (driven by the strong suppression of primordial $\psi(2S)$’s) the ratio essentially saturates and turns out to be consistent with the experimental findings reported by the ALICE collaboration ALICE:2022jeh. Our predictions are significantly larger than the results from the statistical hadronization model Andronic:2019wva; Andronic:2017pug, which level off at approximately $0.05$ in central collisions. As discussed above, the reason for this is that in our transport approach the regeneration of $J/\psi$’s does not reach its equilibrium limit, while the $\psi(2S)$ does, albeit at lower temperatures, recall Figure 7.
@@
Next, we return to scrutinizing the impact of the gluo-dissociation processes on our results by incorporating the pertinent rates into the rate equation [^1] Strictly speaking, this is not following the philosophy of our approach where we adjust the main parameters, i.e., the effective coupling constant, $\alpha_{s}$, in the quasifree rates as well as the thermal relaxation rate of charm quarks, to match SPS and RHIC data Grandchamp:2002wp; Zhao:2010nk. However, the small impact of gluo-dissociation on our results renders this exercise rather obsolete.. In Figure 10, we present results for the time evolution in central collisions (left panel; for direct production) and the centrality dependence (right panel; for inclusive production) of the regenerated, primordial and total $J/\psi$ $R_{\rm AA}$ with and without gluo-dissociation at forward rapidity. Note that a higher rate implies both stronger suppression and increased regeneration. For peripheral and semi-central collisions these two effects essentially compensate each other, while for central collisions, the total $J/\psi$ yield increases by $\sim$6% due to the enhanced regeneration.
@@
Figure 10: The time evolution (left panel) and centrality dependence (right panel, additionally including $B$ feeddown) of the regenerated (red), primordial (orange), and total (blue) $J/\psi$ production, with the bands illustrating the uncertainty as to whether gluo-dissociation rates are accounted for or not. The same values for shadowing and the $c\bar{c}$ cross section as in Figure 7 are used, and the ALICE data are from Ref. ALICE:2023gco.
@@
Let us also briefly come back to the effect of the updated binding energies as compared to our previous calculations, as mentioned in the first paragraph of Sec. $2.$ The impact of the somewhat larger binding energies (most notably around $T\simeq$220 MeV) on regeneration is negligible, as one finds an almost complete compensation between the reduced rates and the larger equilibrium limits. The smaller rates do, however, affect the primordial suppression. The maximal effect for central Pb-Pb collisions amounts to an increase of the primordial contribution by about 0.03 units in terms of the $R_{\rm AA}$.